# A Measurement of the Largest-Scale CMB $E$-mode Polarization with CLASS

### v1

@@ Abstract
We present measurements of large-scale cosmic microwave background (CMB) $E$-mode polarization from the Cosmology Large Angular Scale Surveyor (CLASS) 90 GHz data. Using 115 det-yr of observations collected through 2024 with a variable-delay polarization modulator, we achieved a polarization sensitivity of $78\,\mathrm{\mu K\,arcmin}$, comparable to Planck at similar frequencies (100 and 143 GHz). The analysis demonstrates effective mitigation of systematic errors and addresses challenges to large-angular-scale power recovery posed by time-domain filtering in maximum-likelihood map-making. A novel implementation of the pixel-space transfer matrix is introduced, which enables efficient filtering simulations and bias correction in the power spectrum using the quadratic cross-spectrum estimator. Overall, we achieved an unbiased time-domain filtering correction to recover the largest angular scale polarization, with the only power deficit, arising from map-making non-linearity, being characterized as less than 3%. Through cross-correlation with Planck, we detected the cosmic reionization at $99.4\%$ significance and measured the reionization optical depth $\tau=0.053^{+0.018}_{-0.019}$, marking the first ground-based attempt at such a measurement. At intermediate angular scales ($\ell>30$), our results, both independently and in cross-correlation with Planck, remain fully consistent with Planck’s measurements.
@@ 1 Introduction
The reionization period affects CMB radiation through scattering by free electrons. On the largest scales, the $E$-mode polarization is enhanced by a factor approximately proportional to $\tau^{2}$ (Zaldarriaga, 1997) where $\tau$ is the reionization optical depth. At smaller scales, this scattering suppresses the anisotropy spectrum by a factor of $e^{-2\tau}$ and would also leave detectable imprints in the temperature field if reionization happened inhomogeneously (Gruzinov & Hu, 1998; Dvorkin & Smith, 2009; Namikawa et al., 2021; Raghunathan et al., 2024). To the leading order, a measurement of $\tau$ can constrain the reionization epoch given assumptions of the phase transition (Shull & Venkatesan, 2008; Lewis, 2008). More details of the reionization process can be reconstructed from the full shape of large-scale $E$-mode spectrum (Mortonson & Hu, 2008b; Watts et al., 2020), which can be further complemented by neutral hydrogen probes and direct observations of the ionizing sources to paint a complete picture of the reionization history (Muñoz et al., 2024; Paoletti et al., 2024). Because reionization also bolsters the $B$-mode spectrum, accurately modeling reionization through $E$-mode measurements is crucial for future characterizations of inflationary $B$ modes on the largest scales (Mortonson & Hu, 2008a; Jiang & Namikawa, 2024).
@@ 2 CLASS and the 90 GHz Data Set
Figure 1: CLASS daily observation efficiencies of the 90 GHz telescope since the initial deployment in 2018 through the recent modulator replacement in June 2024. The gray region represents the total detector uptime when the modulator (VPM) was also operational. The yellow region indicates the data volume initially selected at the DataPkg level while the green region shows the fraction retained after the TOD-level selection (see Section 2.3). The telescope “Era” shown in the blue band at the top denotes major changes in the instrument configuration. Gray hatched regions correspond to periods when a plastic closeout was installed. The closeout was first removed on 2021-09-27 but was occasionally reinstalled to protect the instrument from adverse weather conditions. The data with the closeout installed are conservatively left out for the current analysis. Critical changes in the instrument configuration are denoted by vertical bars, with annotations positioned in the direction of their corresponding descriptions.
@@ 2.2 Survey Timeline
The first of the two 90 GHz telescope arrays was deployed in June 2018 and has been observing since then. This telescope was installed on the same mount as the 40 GHz telescope, which began operation in 2016 (Li et al., 2023b, hereafter, L23). Therefore, following the conventions in previous CLASS literature, we denote the first 90 GHz observing season since 2018 as “Era 2”. In response to stability issues in the TES bias range and low optical efficiency (Dahal et al., 2022), the four (out of seven) detector modules with the worst performance were replaced in August 2022. These newly designed modules incorporated updated electromagnetic and electro-thermal circuit modifications, resulting in an overall detector efficiency exceeding 90% (Núñez et al., 2024)—the observation period since the focal plane upgrade is denoted as “Era 3”. Era-3 operations concluded in June 2024 to prepare for the installation of a reflective half-wave plate (RHWP) for the 90 GHz telescope (Eimer et al., 2022; Shi et al., 2024b). This paper focuses only on the 90 GHz data from the first 90 GHz telescope in Era 2 and Era 3, which used the VPM as the modulator.
The timeline and observation efficiency of the 90 GHz observation is summarized in Figure 1, with important events and configuration changes annotated and discussed below:
@@ 2.3 Data Selection
In the first step, 10-minute data segments, each called a DataPkg (Petroff et al., 2020a), were selected based on observation conditions and instrument status. In contrast to the 40 GHz selection criteria, we did not include any daytime data and restricted the Sun position to be below the horizon. We also discarded data acquired while the plastic closeout was installed. [^2] Although this systematic can be removed through filtering for the 40 GHz data (L23), we encountered additional difficulties in modeling the $-\ell$ noise under our current mitigation strategy, yielding only marginal improvements to $-\ell$ sensitivity when including this data set. Nevertheless, it remains unclear whether this limitation stems solely from the closeout since we observed $-\ell$ performance improvement over time that did not align exactly with the closeout status. This resulted in nearly half of the survey data being removed, including all data prior to 2021-09 along with a substantial fraction collected thereafter. The discarded data set could still be used to improve $-\ell$ sensitivity and may be recoverable for future $-\ell$ studies with revised analysis methods. The selected DataPkgs at this stage are shown as the yellow shaded region in Figure 1.
@@ 3.3 Sky Maps
After filtering, the demodulated linear polarization data were weighted and projected onto the sky using a maximum-likelihood map-making algorithm (Dünner et al. 2013; Romero et al. 2020; L23). The core of this method involves constructing noise covariance matrices for data segments that capture most of the temporal and focal plane correlations. For the 90 GHz data, a noise model was estimated approximately every two hours, balancing noise stability, computational cost, and model accuracy. Given the compact form of the noise model, the map solution was obtained using the conjugate gradient method. A challenge arises from the need to derive the noise model from the data themselves, which can introduce signal bias and disrupt linearity, especially when the time-ordered data are less noise dominated. This was mitigated through an iterative template subtraction process, where the map solution from each iteration was subtracted from the time-ordered data before re-estimating the noise model for the next run (L23; Dünner et al. 2013). For the CLASS 90 GHz data, five template iterations were performed. This method has been shown to largely remove bias and restore linearity in map-making across nearly all scales for the 40 GHz data (L23). A detailed characterization of its impact on the 90 GHz analysis is provided in Section 3.4.4.
The CLASS 90 GHz linear polarization maps are shown in Figure 3 in Celestial coordinates. A map of the linear polarization integration time (the hits map) is shown in the last panel of Figure 2. The polarization map depth is $78\,\mathrm{\mu K\,arcmin}$, based on the angular power spectrum in the range $250100<\ell<$ 250 where it is mostly flat.
@@ 3.4.2 Pixel-space Transfer Matrix
To accurately characterize the filtering effect at the map level and the power spectra level, we aim to build a transfer matrix to forward model the filtering in pixel space (BICEP2 Collaboration et al., 2016). A similar approach has been outlined for the CLASS 40 GHz data in L23, but here we refine the matrix estimation and provide more details regarding the assumptions made for this approximation. Since the estimation is computationally prohibitive at the full resolution of the maps, and only the $-\ell$ modes are of interest, we only constructed the matrix for maps at $N_{\mathrm{side}}=16$. To convert the full-resolution maps into lower resolution, we introduce a downgrade operator $\mathcal{D}_{N_{\mathrm{side}}}$ to rebin maps onto $N_{\mathrm{side}}$ grid. The map-level operator ud_grade from Healpix is deliberately avoided to prevent aliasing effects and polarization bias (Sullivan et al., 2024). The downgrade operator $\mathcal{D}_{N_{\mathrm{side}}}$ first transforms the map into spherical harmonic space and then applies a cosine anti-aliasing filter (Benabed et al., 2009; Planck Collaboration Int. XLVI, 2016):
@@
The transfer matrix serves two purposes in this analysis: 1) to forward model the filtering effect, which is useful for processing external template maps and their noise covariance matrices for foreground reduction, and 2) the matrix can be used directly in a quadratic power spectrum estimator to correct for filtering bias in the in a near-optimal way. The implementation details of the latter are discussed in Appendix A. In practice, in order to apply the matrix to an input map, the input needs to be first downgraded to $N_{\mathrm{side}}=N_{F}$. This would cause a double counting of the anti-aliasing filter ($f_{\ell}$, Equation 2) if $N_{F}=16$ because the filter matrix includes $\mathcal{D}_{16}$ by construction (Equation 5). To avoid this, we also evaluate the transfer matrix for $N_{F}=32$ so the extra $f_{\ell}$ at $N_{\mathrm{side}}=32$ has no impact for $\ell\leq 32$. Since covariance matrices at $N_{\mathrm{side}}>16$ resolution for external products are not available, we use $\mathbf{F}_{16,16}$ for noise covariance matrices filtering but apply a window function correction in the harmonic space.
@@ 4.1 Internal Consistency Test
Table 1: Probability-to-exceed (PTE) values for the $-\ell$ and $-\ell$ null tests. The PTE values are tabulated for every null test split and each of the three linear polarization spectra for the two $\ell$-ranges. The cell color indicates the PTE value—red (purple) for PTE values less (greater) than 0.5, meaning that the simulation shows less (more) scatter than the data. Color saturation scales with the extremeness of the PTE (deviation from $0.5$). A subscript third significant digit is included in cells otherwise rounding to 1 or 0. The last row lists the one-tailed p-value from a Kolmogorov Smirnov (KS) test assessing the uniformity of PTE values across splits.
$2\leq\ell\leq 30$ $30131\leq\ell<$ 301
Splits $EE$ $BB$ $EB$ $EE$ $BB$ $EB$
  top/bot   0.11   0.16   0.39   0.07   0.06   0.77
  left/right   0.29   0.76   0.27   0.19   0.94   0.57
  radial   0.34   0.67   0.44   0.88   0.93   0.05
  horizontal   0.63   0.21   0.65   0.59   0.40   0.81
  vertical   0.41   0.97   0.73   0.60   0.52   0.22
  quadrupole   0.97   0.98   0.47   0.54   0.09   0.25
  MUX halves   0.17   0.44   0.75   0.86   0.82   0.91
  MUX parity   0.08   0.88   0.36   0.58   0.67   0.06
  VPM sync.   0.000   0.45   0.11   0.73   0.01   0.84
  bs in/out   0.97   0.72   0.86   0.68   0.96   0.31
  bs pos/neg   0.13   0.29   0.73   0.88   0.12   0.01
  az-east/west   0.23   0.002   0.64   0.03   0.33   0.92
  az-$2\pi$   0.83   0.97   0.99   0.44   0.40   0.93
  az-$4\pi$   0.74   0.53   0.99   0.07   0.68   0.88
  az-velocity   0.43   0.84   $0.99_{7}$   0.39   0.16   0.62
  6h in/out   0.57   0.22   0.34   0.98   0.34   0.03
  moon   0.46   0.24   0.35   0.93   0.88   0.58
  midnight   0.75   0.15   0.30   0.10   0.17   0.02
  survey   0.17   0.02   0.42   0.77   0.87   0.67
KS test 0.29 0.47 0.84 0.61 0.39
@@ 4.2 Consistency with Planck
To isolate the CMB signal, we used the SEVEM CMB component map from the Planck PR4 data release (Planck Collaboration Int. LVII, 2020) to cross-correlate with the CLASS 90 GHz map. The cross-correlation was performed using PolSpice with the same mask used for the null test (Figure 2) that has 70% retained sky fraction after Galactic avoidance. The estimated power spectra have been corrected for the beam and pixel window functions of both maps and are further corrected for the demodulation low-pass filter for the CLASS map. Finally, the agreement with the model is checked by the following $\chi^{2}$ statistic:
@@
where $\hat{C}_{b}/C_{b}$ denote band power from the estimated and fiducial spectra (Planck Collaboration VI, 2020) defined between $30131\leq\ell<$ 301 with $\Delta\ell=10$ bins, $\eta$ is the absolute calibration factor, [^6] The calibration factor correction to the noise covariance matrix is appropriate when the CLASS $E$-mode power spectrum is noise dominated. This procedure has been tested to give unbiased estimation within uncertainties. $F^{1/2}_{b}$ is the CLASS mapping transfer function for cross spectra, [^7] Here we took the square root of the auto-spectra transfer function presented in Figure 5. Strictly speaking, this is not necessarily the same as the transfer function for the cross-spectra between CLASS and Planck due to filtering anisotropy and different sky masks over which the spectra are taken. However, we found negligible differences. and $\mathbf{\Sigma}$ is the band-power covariance matrix. The estimated band power $\hat{C}_{b}$ has also been corrected for the beam and pixel window functions. The $E$-mode transfer function of the Planck map from the PR4 processing (Planck Collaboration Int. LVII, 2020) is negligible at the scales of interest. Here, the covariance matrix was estimated from simulated spectra by pairing 500 CLASS and SEVEM noise simulations with random CMB realizations drawn from the fiducial $\Lambda$CDM model. The covariance matrix was further corrected by the Hartlap factor (Hartlap et al., 2007) of $1.12$ to account for the finite number of simulations. A minimization of the $\chi^{2}$ with respect to $\eta$ gave the best-fit calibration factor $\eta=1.053\pm 0.014$. The fit is driven by the $EE$ spectrum due to the high sample variance in the Planck temperature map. The calibration-adjusted spectra have a $\chi^{2}=53.89$ with respect to the fiducial CMB spectra for $54$ degrees of freedom, or an equivalent PTE value $0.48$. Figure 9 shows the adjusted spectra, along with the residual differences expressed in $\chi$. Of note, the absolute calibration leverages the CMB power spectra at scales $\ell>30$, which is proportional to $A_{\mathrm{s}}e^{-2\tau}$. Therefore, it does not compromise the experiment’s ability to measure $\tau$ independently of the calibrator at large angular scales ($\ell<30$), to the extent that $A_{\mathrm{s}}e^{-2\tau}$ is considered a well-constrained parameter.

### v2

@@ Abstract
We present measurements of large-scale cosmic microwave background (CMB) $E$-mode polarization from the Cosmology Large Angular Scale Surveyor (CLASS) 90 GHz data. Using 115 det-yr of observations collected through 2024 with a variable-delay polarization modulator, we achieved a polarization sensitivity of $82\,\mathrm{\mu K\,arcmin}$, comparable to Planck at similar frequencies (100 and 143 GHz). The analysis demonstrates effective mitigation of systematic errors and addresses challenges to large-angular-scale power recovery posed by time-domain filtering in maximum-likelihood map-making. A novel implementation of the pixel-space transfer matrix is introduced, which enables efficient filtering simulations and bias correction in the power spectrum using the quadratic cross-spectrum estimator. Overall, we achieved an unbiased time-domain filtering correction to recover the largest angular scale polarization, with the only power deficit, arising from map-making non-linearity, being characterized as less than 3%. Through cross-correlation with Planck, we detected the cosmic reionization at $99.4\%$ significance and measured the reionization optical depth $\tau=0.053^{+0.018}_{-0.019}$, marking the first ground-based attempt at such a measurement. At intermediate angular scales ($\ell>30$), our results, both independently and in cross-correlation with Planck, remain fully consistent with Planck’s measurements.
@@ 1 Introduction
The reionization period affects CMB radiation through scattering by free electrons. On the largest scales, the $E$-mode polarization is enhanced by a factor approximately proportional to $\tau^{2}$ (Zaldarriaga, 1997) where $\tau$ is the reionization optical depth. At smaller scales, this scattering suppresses the anisotropy spectrum by a factor of $e^{-2\tau}$ and would also leave detectable imprints in the temperature field if reionization happened inhomogeneously (Gruzinov & Hu, 1998; Dvorkin & Smith, 2009; Namikawa et al., 2021; Raghunathan et al., 2024). To the leading order, a measurement of $\tau$ can constrain the reionization epoch given assumptions of the phase transition (Shull & Venkatesan, 2008; Lewis, 2008). More details of the reionization process can be reconstructed from the full shape of large-scale $E$-mode spectrum (Mortonson & Hu, 2008b; Watts et al., 2020), which can be further complemented by neutral hydrogen probes and direct observations of the ionizing sources to paint a complete picture of the reionization history (Muñoz et al., 2024; Paoletti et al., 2025). Because reionization also bolsters the $B$-mode spectrum, accurately modeling reionization through $E$-mode measurements is crucial for future characterizations of inflationary $B$ modes on the largest scales (Mortonson & Hu, 2008a; Jiang & Namikawa, 2024).
@@ 2 CLASS and the 90 GHz Data Set
Figure 1: CLASS daily observation efficiencies (see definition in text) of the 90 GHz telescope since the initial deployment in 2018 through the recent modulator replacement in June 2024. The gray region represents the total detector uptime when the modulator (VPM) was also operational. The yellow region indicates the data volume initially selected at the DataPkg level while the green region shows the fraction retained after the TOD-level selection (see Section 2.3). The telescope “Era” shown in the blue band at the top denotes major changes in the instrument configuration. Gray hatched regions correspond to periods when a plastic closeout was installed. The closeout was first removed on 2021-09-27 but was occasionally reinstalled to protect the instrument from adverse weather conditions. The data with the closeout installed are conservatively left out for the current analysis. Critical changes in the instrument configuration are denoted by vertical bars, with annotations positioned in the direction of their corresponding descriptions.
@@ 2.2 Survey Timeline
The first of the two 90 GHz telescope arrays was deployed in June 2018 and has been observing since then. This telescope was installed on the same mount as the 40 GHz telescope, which began operation in 2016 (Li et al., 2023b, hereafter, L23). Therefore, following the conventions in previous CLASS literature, we denote the first 90 GHz observing season since 2018 as “Era 2”. In response to stability issues in the TES bias range and low optical efficiency (Dahal et al., 2022), the four (out of seven) detector modules with the worst performance were replaced in August 2022. These newly designed modules incorporated updated electromagnetic and electro-thermal circuit modifications, resulting in an overall detector optical efficiency exceeding 90% (Núñez et al., 2024)—the observation period since the focal plane upgrade is denoted as “Era 3”. Era-3 operations concluded in June 2024 to prepare for the installation of a reflective half-wave plate (RHWP) for the 90 GHz telescope (Eimer et al., 2022; Shi et al., 2024b). This paper focuses only on the 90 GHz data from the first 90 GHz telescope in Era 2 and Era 3, which used the VPM as the modulator.
The timeline and observation efficiency (defined as the ratio between the achieved integration time and a 24-hour observation with a fully populated focal plane) of the 90 GHz observation is summarized in Figure 1, with important events and configuration changes annotated and discussed below:
@@ 2.3 Data Selection
In the first step, 10-minute data segments, each called a DataPkg (Petroff et al., 2020a), were selected based on observation conditions and instrument status. In contrast to the 40 GHz selection criteria, we did not include any daytime data and restricted the Sun position to be below the horizon. We also discarded data acquired while the plastic closeout was installed. [^2] Although this systematic can be removed through filtering for the 40 GHz data (L23), we encountered additional difficulties in modeling the $-\ell$ noise under our current mitigation strategy, yielding only marginal improvements to $-\ell$ sensitivity when including this data set. Nevertheless, it remains unclear whether this limitation stems solely from the closeout since we observed $-\ell$ performance improvement over time that did not align exactly with the closeout status. This resulted in nearly half of the survey data being removed, including all data prior to 2021-09 along with a substantial fraction collected thereafter. The discarded data set could still be used to improve $-\ell$ sensitivity and may be recoverable for future $-\ell$ studies with revised analysis methods. The selected DataPkgs at this stage, amounting to 190 det-yr, are shown as the yellow shaded region in Figure 1.
@@ 3.3 Sky Maps
After filtering, the demodulated linear polarization data were weighted and projected onto the sky using a maximum-likelihood map-making algorithm (Dünner et al. 2013; Romero et al. 2020; L23). The core of this method involves constructing noise covariance matrices for data segments that capture most of the temporal and focal plane correlations. For the 90 GHz data, a noise model was estimated approximately every two hours, balancing noise stability, computational cost, and model accuracy. Given the compact form of the noise model, the map solution was obtained using the conjugate gradient method. A challenge arises from the need to derive the noise model from the data themselves, which can introduce signal bias and disrupt linearity, especially when the time-ordered data are less noise-dominated. This was mitigated through an iterative template subtraction process, where the map solution from each iteration was subtracted from the time-ordered data before re-estimating the noise model for the next run (L23; Dünner et al. 2013). For the CLASS 90 GHz data, five template iterations were performed. This method has been shown to largely remove bias and restore linearity in map-making across nearly all scales for the 40 GHz data (L23). A detailed characterization of its impact on the 90 GHz analysis is provided in Section 3.4.4.
The CLASS 90 GHz linear polarization maps are shown in Figure 3 in Celestial coordinates. A map of the linear polarization integration time (the hits map) is shown in the last panel of Figure 2. The polarization map depth is $82\,\mathrm{\mu K\,arcmin}$, based on the angular power spectrum in the range $250100<\ell<$ 250 where it is mostly flat.
@@ 3.4.2 Pixel-space Transfer Matrix
To accurately characterize the filtering effect at the map level and the power spectra level, we aim to build a transfer matrix to forward model the filtering in pixel space (BICEP2 Collaboration et al., 2016). A similar approach has been outlined for the CLASS 40 GHz data in L23, but here we refine the matrix estimation and provide more details regarding the assumptions made for this approximation. Since the estimation is computationally prohibitive at the full resolution of the maps, and only the $-\ell$ modes are of interest, we only constructed the matrix for maps at $N_{\mathrm{side}}=16$ (the pixel number scales as $N_{\mathrm{side}}^{2}$). To convert the full-resolution maps into lower resolution, we introduce a downgrade operator $\mathcal{D}_{N_{\mathrm{side}}}$ to rebin maps onto $N_{\mathrm{side}}$ grid. The map-level operator ud_grade from Healpix is deliberately avoided to prevent aliasing effects and polarization bias (Sullivan et al., 2025). The downgrade operator $\mathcal{D}_{N_{\mathrm{side}}}$ first transforms the map into spherical harmonic space and then applies a cosine anti-aliasing filter (Benabed et al., 2009; Planck Collaboration Int. XLVI, 2016):
@@
The transfer matrix serves two purposes in this analysis: 1) to forward model the filtering effect, which is useful for processing external template maps and their noise covariance matrices for foreground reduction, and 2) the matrix can be used directly in a quadratic power spectrum estimator to correct for filtering bias in a near-optimal way. The implementation details of the latter are discussed in Appendix A. In practice, in order to apply the matrix to an input map, the map must first be downgraded (with $f_{\ell}$ from Equation 2) to $N_{\mathrm{side}}=N_{F}$ to match the matrix’s dimensions. This would cause a double counting of the anti-aliasing filter, as the filter matrix inherently includes $\mathcal{D}_{16}$ (Equation 5). To avoid this, we evaluate the transfer matrix at $N_{F}=32$, ensuring that the additional $f_{\ell}$ for $N_{\mathrm{side}}=32$ has no effect within the target range $\ell\leq 32$ due to the filter design (Equation 2). Since covariance matrices at $N_{\mathrm{side}}>16$ resolution for external products are not available, we use $\mathbf{F}_{16,16}$ for noise covariance matrices filtering but apply a window function correction in the harmonic space.
@@ 4.1 Internal Consistency Test
Figure 9: Probability-to-exceed (PTE) values for the $-\ell$ and $-\ell$ null tests. The PTE values are tabulated for every null test split and each of the three linear polarization spectra for the two $\ell$-ranges. The cell color indicates the PTE value—red (purple) for PTE values less (greater) than 0.5, meaning that the simulation shows less (more) scatter than the data. Color saturation scales with the extremeness of the PTE (deviation from $0.5$). A subscript third significant digit is included in cells otherwise rounding to 1 or 0. The last row lists the one-tailed p-value from a Kolmogorov Smirnov (KS) test assessing the uniformity of PTE values across splits.
@@ 4.2 Consistency with Planck
To isolate the CMB signal, we used the SEVEM CMB component map from the Planck PR4  [^6] We used Planck PR4 mainly for its improved systematic error control on large angular scales in polarization. data release (Planck Collaboration Int. LVII, 2020) to cross-correlate with the CLASS 90 GHz map. The cross-correlation was performed using PolSpice with the same mask used for the null test (Figure 2) that has 70% retained sky fraction after Galactic avoidance. The estimated power spectra have been corrected for the beam and pixel window functions of both maps and are further corrected for the demodulation low-pass filter for the CLASS map. Finally, the agreement with the model is checked by the following $\chi^{2}$ statistic:
@@
where $\hat{C}_{b}/C_{b}$ denote band power from the estimated and fiducial spectra (Planck Collaboration VI, 2020) defined between $30131\leq\ell<$ 301 with $\Delta\ell=10$ bins, $\eta$ is the absolute calibration factor, [^7] The calibration factor correction to the noise covariance matrix is appropriate when the CLASS $E$-mode power spectrum is noise dominated. This procedure has been tested to give unbiased estimation within uncertainties. $F^{1/2}_{b}$ is the CLASS mapping transfer function for cross spectra, [^8] Here we took the square root of the auto-spectra transfer function presented in Figure 5. Strictly speaking, this is not necessarily the same as the transfer function for the cross-spectra between CLASS and Planck due to filtering anisotropy and different sky masks over which the spectra are taken. However, we found negligible differences. and $\mathbf{\Sigma}$ is the band-power covariance matrix. The estimated band power $\hat{C}_{b}$ has also been corrected for the beam and pixel window functions. The $E$-mode transfer function of the Planck map from the PR4 processing (Planck Collaboration Int. LVII, 2020) is negligible at the scales of interest. Here, the covariance matrix was estimated from simulated spectra by pairing 500 CLASS and SEVEM noise simulations with random CMB realizations drawn from the fiducial $\Lambda$CDM model. The covariance matrix was further corrected by the Hartlap factor (Hartlap et al., 2007) of $1.12$ to account for the finite number of simulations. A minimization of the $\chi^{2}$ with respect to $\eta$ gave the best-fit calibration factor $\eta=1.053\pm 0.014$. The fit is driven by the $EE$ spectrum due to the higher sample variance of the $TE$ spectrum from the Planck temperature map. The calibration-adjusted spectra have a $\chi^{2}=53.89$ with respect to the fiducial CMB spectra for $54$ degrees of freedom, or an equivalent PTE value $0.48$. Figure 10 shows the adjusted spectra, along with the residual differences expressed in $\chi$. Of note, the absolute calibration leverages the CMB power spectra at scales $\ell>30$, which is proportional to $A_{\mathrm{s}}e^{-2\tau}$. Therefore, it does not compromise the experiment’s ability to measure $\tau$ independently of the calibrator at large angular scales ($\ell<30$), to the extent that $A_{\mathrm{s}}e^{-2\tau}$ is considered a well-constrained parameter.